OpenStudy (anonymous):

Select the proper equation in point-slope form for the given point and slope. Point: (1, –7); Slope: y + 7 = (x – 1) y –7 = (x + 1) y + 7 = (x + 1) y –7 = (x – 1)

OpenStudy (whpalmer4):

Okay, I just showed you how to do this, time to do it!

OpenStudy (anonymous):

1. but do you know why? i also answered a question like this all you need to do is use the equation.

OpenStudy (anonymous):

y- y1= m(x-x1) the point is 1,-7 so y- (-7) = x-1 so its y+7= x-1

OpenStudy (anonymous):

@whpalmer4 Hey so what? It’s a practice thingy, no points for me

OpenStudy (whpalmer4):

So practice!

OpenStudy (anonymous):

its not mandatory

OpenStudy (whpalmer4):

It is if you want to be able to do this stuff. Hint: you'll be seeing a lot of it.

OpenStudy (anonymous):

ik my bro is telling me